OpenStudy (anonymous):

find the exact value of csc(arctan 11/60)

OpenStudy (anonymous):

Let \(\theta=\arctan\dfrac{11}{60}\), so that \(\tan\theta=\dfrac{11}{60}\). Consider the triangle,  Find the missing side, then find the cosecant of the angle \(\theta\).